OpenStudy (anonymous):

1/2t + 6/t =1 1/2t + 12/2t =13/2t 13/2t = 1 divide both by 2 to leave x alone & its 13/2

OpenStudy (anonymous):

is this right ?

OpenStudy (anonymous):

x=13/2

OpenStudy (anonymous):

what does the question ask you to do?

OpenStudy (anonymous):

just to solve for t

OpenStudy (anonymous):

1/2t + 6/t =1

OpenStudy (anonymous):

yes, your answer is correct

OpenStudy (anonymous):

it is 13/2 which is 6.5

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

the easiest way to do this is to multiply both sides by t leave you with 1/2+6=t Or, t = 6+1/2 or 6.5 or 13/2